# The Hessian of tall-skinny networks is easy to invert

### v1

@@ 1 Introduction
The Hessian of a deep net is the matrix of second-order mixed partial derivatives of its loss with respect to its parameters. Decades ago, when deep nets had only hundreds or thousands of parameters, the Hessian matrix could be inverted to implement optimizers that converged much faster than gradient descent [Watrous1987, Barnard1992]. But for large modern deep nets, using the Hessian has become less practical: The Hessian of a model with a billion parameters would have a quintillion entries, which is far larger than can be stored, multiplied, factorized, or inverted, even in the largest data centers. A common workaround is to approximate the Hessian as a low-rank matrix [Webb1988, lbfgs-deep-net] or as a diagonal matrix [becker-lecun-89, adagrad, adam]. Such approximations make it easier to apply the inverse of the Hessian to a vector. This article shows how to compute and apply the inverse of the Hessian exactly without storing the Hessian and without storing its inverse. The Hessian-inverse-vector product can be computed in time and storage that scales linearly with the number of layers in the model, and cubically in the number of parameters and activations in each layer.
Pearlmutter [pearlmutter94] showed the product of the Hessian with a fixed vector (the so-called Hessian-vector product) can be computed in time and storage that scales linearly with the number of layers in the network. The scaling behavior of Pearlmutter’s method is much better than the cubic scaling of the naive algorithm that first computes the Hessian matrix and then multiplies it by the vector. The method transforms the original network into a new network whose gradient is the desired Hessian-vector product. To compute the Hessian-vector product, one then just applies backpropagation to the new network. Given a way to compute the Hessian-vector product, one can indirectly compute the Hessian-inverse-vector product via, say Krylov iterations like Conjugate Gradient. However, the quality of the result would then depend on the conditioning of the Hessian, which is notoriously poor for deep nets [ying-behrooz-hessian-spectrum]. Unfortunately, there seems to exist no variant of Pearlmutter’s trick to compute the Hessian-inverse-vector products directly.
@@ 2 Overview
Algorithm 1 Compute the hessian-inverse-vector product by solving $(H+\epsilon I)x=b$
@@ 5 The Hessian
To solve systems of the form $Hx=b$, one could use Krylov methods to repeatedly multiply by $H$ without forming $H$ (a possibility Pearlmutter considered [pearlmutter94]). This would require applying $H$ some number of times that depends on the condition number of $H$. However, the next section shows how to solve systems of the form $H^{-1}x=b$ with only $\max(a,p)$ times more operations than are needed to compute $Hx$.
@@ 6 Applying the inverse of the Hessian
The pivoting we’ll apply reorders the blocks of the variables $x,y,z$ from $x_{1},\ldots,x_{L},y_{1},\ldots,y_{L},z_{1},\ldots,z_{L}$ to $x_{1},y_{1},z_{1},\dots,x_{L},y_{L},z_{L}$. This pivoting operation acts as a kind of transpose operation on block matrices (its generalization is called a “commutation matrix” in [magnus-neudecker]). When applied to $\mathcal{K}$, it turns it into a block-tri-diagonal matrix. To get a better feel for this operation, denote the $ij$th block of $\mathcal{K}$ by $\mathcal{K}_{ij}$ and the $uv$th sub-block of this block by $\mathcal{K}_{ij,uv}$. Similarly, $x$ denote a vector conformant with $\mathcal{K}$, with $x_{j}$ denoting the column vector that multiplies each block $\mathcal{K}_{\cdot j}$ and denote the $v$th sub-block of $x_{j}$ by $x_{jv}$, which multiplies $\mathcal{K}_{\cdot j,\cdot v}$. We say that $x$ is $j$-major because as we traverse the entries of $x$ from top to bottom, the index $j$ increments more slowly than the $v$ index.
@@ Appendix A The Hessian
as $dx\to 0$. This convention readily implies the Hessian of a vector-valued function: If $g:\mathop{\mathbb{R}}^{d}\to\mathop{\mathbb{R}}^{o}$, then $\nabla_{xx}g(x)\in\mathop{\mathbb{R}}^{o\times d^{2}}$ is the unique matrix such that $\mathrm{vec}\left(\nabla_{x}g(x+dx)-\nabla_{x}g(x)\right)\to\nabla_{xx}g(x)\;dx$. This convention also readily accommodates the chain rule. For example, the gradient of $h(x)\equiv f(g(x))$ for matrix-valued $f$ and $g$ can be written as $\nabla f\nabla g$ as expected. It also implies partial derivatives like $\nabla_{yz}g$ for $g:\mathop{\mathbb{R}}^{|x|}\to\mathop{\mathbb{R}}^{|g|}$. If $y\in\mathop{\mathbb{R}}^{|y|}$ and $z\in\mathop{\mathbb{R}}^{|z|}$ are restrictions of $x\in\mathop{\mathbb{R}}^{|x|}$ to some $|y|$ and $|z|$-dimensional subsets, then $\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g|\times|z|}$, and $\nabla_{yz}g(x)=\nabla_{y}\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g||z|\times|y|}$. See Chapter 6 of Magnus & Neudecker [magnus-neudecker] for a deeper treatment of higher order derivatives of vector-valued functions.
@@ Appendix B Equation (9) is Pearlmutter’s Hessian-vector multiplication algorithm
We showed that Equation (9) makes it possible to compute Hessian-vector product $Hv$ in time linear in $L$. These operations are equivalent to Pearlmutter’s [pearlmutter94] algorithm, a framework to compute Hessian-vector products in networks with arbitrary topologies. This section specializes Pearlmutter’s machinery to the pipeline topology, and shows that the operations it produces coincide exactly with those of Equation (9).

### v2

@@ 1 Introduction
The Hessian of a deep net is the matrix of second-order mixed partial derivatives of its loss with respect to its parameters. Decades ago, when deep nets had only hundreds or thousands of parameters, the Hessian matrix could be inverted to implement optimizers that converged much faster than gradient descent [9, 1]. But for large modern deep nets, using the Hessian has become less practical: The Hessian of a model with a billion parameters would have a quintillion entries, which is far larger than can be stored, multiplied, factorized, or inverted, even in the largest data centers. A common workaround is to approximate the Hessian as a low-rank matrix [10, 6] or as a diagonal matrix [2, 3, 5]. Such approximations make it easier to apply the inverse of the Hessian to a vector. This article shows how to compute and apply the inverse of the Hessian exactly without storing the Hessian and without storing its inverse. The Hessian-inverse-vector product can be computed in time and storage that scales linearly with the number of layers in the model, and cubically in the number of parameters and activations in each layer.
Pearlmutter [8] showed the product of the Hessian with a fixed vector (the so-called Hessian-vector product) can be computed in time and storage that scales linearly with the number of layers in the network. The scaling behavior of Pearlmutter’s method is much better than the cubic scaling of the naive algorithm that first computes the Hessian matrix and then multiplies it by the vector. The method transforms the original network into a new network whose gradient is the desired Hessian-vector product. To compute the Hessian-vector product, one then just applies backpropagation to the new network. Given a way to compute the Hessian-vector product, one can indirectly compute the Hessian-inverse-vector product via, say Krylov iterations like Conjugate Gradient. However, the quality of the result would then depend on the conditioning of the Hessian, which is notoriously poor for deep nets [4]. Unfortunately, there seems to exist no variant of Pearlmutter’s trick to compute the Hessian-inverse-vector products directly.
@@ 2 Overview
Algorithm 1 Compute the hessian-inverse-vector product by solving $(H+\epsilon I)x=b$.
@@ 5 The Hessian
To solve systems of the form $Hx=b$, one could use Krylov methods to repeatedly multiply by $H$ without forming $H$ (a possibility Pearlmutter considered [8]). This would require applying $H$ some number of times that depends on the condition number of $H$. However, the next section shows how to solve systems of the form $H^{-1}x=b$ with only $\max(a,p)$ times more operations than are needed to compute $Hx$.
@@ 6 Applying the inverse of the Hessian
The pivoting we’ll apply reorders the blocks of the variables $x,y,z$ from $x_{1},\ldots,x_{L},y_{1},\ldots,y_{L},z_{1},\ldots,z_{L}$ to $x_{1},y_{1},z_{1},\dots,x_{L},y_{L},z_{L}$. This pivoting operation acts as a kind of transpose operation on block matrices (its generalization is called a “commutation matrix” in [7]). When applied to $\mathcal{K}$, it turns it into a block-tri-diagonal matrix. To get a better feel for this operation, denote the $ij$th block of $\mathcal{K}$ by $\mathcal{K}_{ij}$ and the $uv$th sub-block of this block by $\mathcal{K}_{ij,uv}$. Similarly, $x$ denote a vector conformant with $\mathcal{K}$, with $x_{j}$ denoting the column vector that multiplies each block $\mathcal{K}_{\cdot j}$ and denote the $v$th sub-block of $x_{j}$ by $x_{jv}$, which multiplies $\mathcal{K}_{\cdot j,\cdot v}$. We say that $x$ is $j$-major because as we traverse the entries of $x$ from top to bottom, the index $j$ increments more slowly than the $v$ index.
@@ Appendix A The Hessian
as $dx\to 0$. This convention readily implies the Hessian of a vector-valued function: If $g:\mathop{\mathbb{R}}^{d}\to\mathop{\mathbb{R}}^{o}$, then $\nabla_{xx}g(x)\in\mathop{\mathbb{R}}^{o\times d^{2}}$ is the unique matrix such that $\mathrm{vec}\left(\nabla_{x}g(x+dx)-\nabla_{x}g(x)\right)\to\nabla_{xx}g(x)\;dx$. This convention also readily accommodates the chain rule. For example, the gradient of $h(x)\equiv f(g(x))$ for matrix-valued $f$ and $g$ can be written as $\nabla f\nabla g$ as expected. It also implies partial derivatives like $\nabla_{yz}g$ for $g:\mathop{\mathbb{R}}^{|x|}\to\mathop{\mathbb{R}}^{|g|}$. If $y\in\mathop{\mathbb{R}}^{|y|}$ and $z\in\mathop{\mathbb{R}}^{|z|}$ are restrictions of $x\in\mathop{\mathbb{R}}^{|x|}$ to some $|y|$ and $|z|$-dimensional subsets, then $\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g|\times|z|}$, and $\nabla_{yz}g(x)=\nabla_{y}\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g||z|\times|y|}$. See Chapter 6 of Magnus & Neudecker [7] for a deeper treatment of higher order derivatives of vector-valued functions.
@@ Appendix B Equation (9) is Pearlmutter’s Hessian-vector multiplication algorithm
We showed that Equation (9) makes it possible to compute Hessian-vector product $Hv$ in time linear in $L$. These operations are equivalent to Pearlmutter’s [8] algorithm, a framework to compute Hessian-vector products in networks with arbitrary topologies. This section specializes Pearlmutter’s machinery to the pipeline topology, and shows that the operations it produces coincide exactly with those of Equation (9).